Shallypeace:

In arithmetic progression the difference between the 8th and 4th term is 20 and the 8th term is ½ times the 4th term. Find the common difference and the first term in the sequence

Timmyspu:

So what do you think is the awnser

Shallypeace:

I don't know

Timmyspu:

I can help you but I cant give you a direct awnser

Shallypeace:

But if anyone knows please show how you solved it

Shallypeace:

OK

Timmyspu:

I cant help you that well but I can try my best

Shallypeace:

OK no problem

Shallypeace:

OK thanks a lot❤

Timmyspu:

No prob

Timmyspu:

do you need any more help

Shallypeace:

No for now

Timmyspu:

Ok just tell me if you need any more help

Shallypeace:

Do you have any other method

Shallypeace:

Like using simultaneous or elimination method

Mercury:

generic arithmetic sequence formula \[a_{n}=a_{1}+(n-1)d\] for the 4th and 8 terms: \[a_{8}=a_{1}+(8-1)d\]\[a_{4}=a_{1}+(4-1)d\] it says "the difference between the two terms is 20" so we can subtract a8 - a4 to get a8 - a4 = (8-1)d - (4-1)d = 20 solve for d to get the common difference "8th term is ½ times the 4th term" so a8 = (1/4)a4 you can plug (1/4)a4 into the a8 equation to get \[(1/4)a_{4}=a_{1}+(8-1)d\] you can multiply this by 4 on both sides to get a4 = 4(a1 + (8-1)d) now, looking at the original a4 equation\[a_{4}=a_{1}+(4-1)d\] you can set these a4 equations equal to each other, plug in the d-value from before, and solve for a1